OpenStudy (anonymous):

What is the 9th term of the geometric sequence where a1 = -5 and a6 = -5,120

hartnn (hartnn):

do you know the formula for n'th term of geometric series ?

hartnn (hartnn):

\(\Large a_n = a_1 r^{n-1}\) plug in a1 = -5 and a6 = -5,120, n = 6

hartnn (hartnn):

and find 'r'

OpenStudy (anonymous):

where doe -5120 go in the formula

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

its the 6th term, so its \(\Large a_6 \) when you plug in n=6 the left side = a_6 = -5120

hartnn (hartnn):

\(\Large -5120 = -5 r^{6-1}\) find r

OpenStudy (anonymous):

-5120=-5r^5 where do i go after that

hartnn (hartnn):

divide both sides by 5

hartnn (hartnn):

i mean -5

OpenStudy (anonymous):

1024=r^5

hartnn (hartnn):

yup now take 5th root on both sides

OpenStudy (anonymous):

how do you do that

hartnn (hartnn):

you can use calculator , right ?

OpenStudy (anonymous):

is it 4

hartnn (hartnn):

yes! r= 4 now use the same formula again to get 9th term this time r= 9

hartnn (hartnn):

\(\Large a_9 = a_1 r^{9-1} \\ \Large a_9 = -5 \times 4^8 = ...\)

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

-327,680

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

What is the sum of the arithmetic sequence 151, 137, 123, …, if there are 26 terms

hartnn (hartnn):

could you find the common difference ?

OpenStudy (anonymous):

its -14

hartnn (hartnn):

thats correct, d= -14 a1 = 1st term = 151 n = number of terms = 26

hartnn (hartnn):

just use this formula directly \(\Large S_n = (n/2) (2a_1 + (n-1)d)\)

OpenStudy (anonymous):

-624?

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

thank you so much

hartnn (hartnn):

welcome ^_^